For the following reaction, what is the percentage yield of magnesium nitride if 1.70 g of it are produced when 1.36 g of Mg reacts with 0.695 g of N2? (10 pts)
3Mg + N2 —> Mg3N2
1
Expert's answer
2020-05-16T13:42:40-0400

3Mg + N2 —> Mg3N2

1. Calculating limiting reactant:

n(Mg) = m(Mg) / [3 × Mr(Mg)] = 1.36 g / (3 × 24 g/mol) = 0.019 mol

n(N2) = m(N2) / Mr(N2) = 0.695 g / 28 g/mol = 0.025 mol

As n(Mg) < n(N2), Mg is a limiting reactant.

2. Calculating the theoretical yield of Mg3N2:

m(Mg3N2) = m(Mg) × Mr(Mg3N2) / [3 × Mr(Mg)] = n(Mg) × Mr(Mg3N2) = 0.019 mol × 101 g/mol = 1.92 g

3. Calculating the percentage yield:

As the theoretical yield equals 1.92 g and actual yield equals 1.70 g, then:

Percentage yield = actual yield / theoretical yield × 100% = 1.70 g / 1.92 g × 100% = 88.5%


Answer: 88.5%